OpenStudy (anonymous):

3sin^3 (x) csc(x) + cos^2 (x) + 2cos (−x) cos (x) = A. tan^2 (x) + 1 B. cos^2 (x) C. 3 D. sin^2 (x)

OpenStudy (anonymous):

I was thinking it was B. but im not too sure. . .

OpenStudy (anonymous):

\[3\sin^3x \frac{ 1 }{ sinx }+\cos^2x+2\cos^x\] \[3\sin^2x+3\cos^2x\] simplify and answer will be...........

OpenStudy (anonymous):

3

OpenStudy (anonymous):

@ercumpian yes

OpenStudy (anonymous):

Thanx for the help @ASAAD123

OpenStudy (anonymous):

welcome ^_|_^ -|-